Solve please:
The entrance to an athletic field is in the shape of a parabolic archway. The archway is modeled by the equation d = 12x = x^2, where d represents the distance, in feet, that the arch is above the ground for any x value.
a. For what values of x will the arch be 20 feet above the ground?
b. How many feet wide is the arch?

Answer by stanbon(75887)

The entrance to an athletic field is in the shape of a parabolic archway. The archway is modeled by the equation d = 12x - x^2, where d represents the distance, in feet, that the arch is above the ground for any x value.
a. For what values of x will the arch be 20 feet above the ground?
Solve::
12x - x^2 = 20
x^2 - 12x + 20 = 0
(x-2)(x-10) = 0
Ans: x = 2 ft and x = 10 ft
-------------------------------------

b. How many feet wide is the arch?
Solve::
12x - x^2 = 0
x^2 - 12x = 0
x(x-12) = 0
x = 0 and x = 12
------
Ans: 12 ft.
----------
Cheers,
Stan H.
--------